### v1

# State Frequency Estimation for Anomaly Detection

Abstract
Many works have studied the efficacy of state machines for detecting anomalies within NetFlows. These works typically learn a model from unlabeled data and compute anomaly scores for arbitrary traces based on their likelihood of occurrence or how well they fit within the model. However, these methods do not dynamically adapt their scores based on the traces seen at test time. This becomes a problem when an adversary produces seemingly common traces in their attack, causing the model to miss the detection by assigning low anomaly scores. We propose SEQUENT, a new approach that uses the state visit frequency to adapt its scoring for anomaly detection dynamically. SEQUENT subsequently uses the scores to generate root causes for anomalies. These allow the grouping of alarms and simplify the analysis of anomalies. Our evaluation of SEQUENT on three NetFlow datasets indicates that our approach outperforms existing methods, demonstrating its effectiveness in detecting anomalies.
I Introduction
NetFlow encapsulates summarized network statistics of a connection between two hosts and can be used for traffic analysis. Studies have shown the effectiveness of state machines in detecting anomalies on NetFlows in an unsupervised manner [24, 10, 12]. Typically, these studies learn a model from benign flows and then compute an anomaly score based on the likelihood of a trace given the model.
Although a standard approach in machine learning (ML), we argue that this has two limitations making it not well-suited for detecting anomalies we see in practice. Firstly, many benign flows occur with a low probability, resulting in a large number of false alarms. Secondly, malicious flows tend to resemble common flows, giving them large probabilities and thus making it hard to raise alarms. In this paper, we propose a way to overcome these limitations by using a key feature of the state machine model: its behavioral structure. Instead of computing likelihoods, we compute visit frequencies of different states within the model. Each state essentially represents a distinct type of network traffic behavior (a cluster or latent value). When we observe behavior much more frequently than expected, we raise an alarm. Intuitively, benign flows with small probability thus do not raise alarms when they occur as frequently as expected. Moreover, malicious flows with high probability do raise alarms when they occur more frequently than expected. Note that this also detects infrequent malicious flows that were not observed at training since their expected count is assumed to be very small. The price to pay is that now normal flows raise alarms when they occur more frequently. Although this does happen in network traffic, we believe and demonstrate in experiments that this is a price worth paying.
We introduce SEQUENT, an approach that learns a state machine from benign NetFlows to model sequential network behavior. The process begins by discretizing the NetFlow features into a symbolic representation. We develop a discretization method [3] inspired by the Glove algorithm [25]. We discretize each selected feature independently and then map each NetFlow to a symbol, which is the concatenation of the individual discretized symbols.
SEQUENT then runs each trace observed at test time on the learned model and computes the set of states visited by the trace together with their corresponding visit frequency. SEQUENT compares the visit frequencies to the expected frequencies observed at train time and raises an alarm whenever a type of behavior occurs more frequently than expected.
A well-known challenge faced by many security analysts in their daily routine is the large volume of alerts raised by the anomaly detection system (ADS) [21, 14, 1], turning the inspection of alerts to a labor-intensive task. The sheer volume of alerts makes it difficult for analysts to prioritize which anomalies to inspect first. SEQUENT addresses this issue by computing root causes (state identifiers) for each detected anomaly, which allows grouping and ranking of anomalies (e.g., using the frequency of a root cause). Additionally, SEQUENT establishes links between the root causes and their corresponding NetFlows. The aim is to provide analysts with concrete instances of anomalous behavior, enabling a quick and in-depth analysis of anomalies.
We empirically demonstrate SEQUENT’s effectiveness in the task of detecting various types of network anomalies using three publicly available NetFlow datasets [9, 11, 17]. Our experiment results demonstrate that SEQUENT is highly effective in the detection of network anomalies, outperforming existing state-of-the-art unsupervised methods. The main contributions of this paper are summarized as:
  • We present SEQUENT, a novel approach that employs state visit frequencies of a state machine for anomaly detection.
  • SEQUENT offers the capability to group and rank anomalies, empowering analysts to filter and prioritize which anomalies they should investigate.
  • We compare SEQUENT against existing state-of-the-art unsupervised methods on three different datasets, demonstrating its effectiveness and generalizability.
II Preliminaries
II-A State Machine
A state machine, formally known as a finite state machine or finite state automaton, is a mathematical model used for modeling the sequential behavior of a system. It is formally defined as a 5-tuple $(\Sigma,Q,q_{0},\delta,F)$[26], where:
  • $\Sigma$ is a finite alphabet
  • $Q$ is a finite set of states
  • $q_{0}\in Q$ is a unique start state
  • $\delta:Q\times\Sigma\to Q$ is the transition function
  • $F\subseteq Q$ is the set of final states
Given an input trace $t$ (a sequence of events, where each event $e\in\Sigma$) and a system $S$, one can use the state machine to determine in which state $S$ would end up, thereby understanding the behavior the system would exhibit. An illustrative example of a state machine is shown in Figure 1. The state machine in this example consists of only one final state, $q_{3}$.
Figure 1: An illustrative example of a state machine. At the end of the input trace, $aabab$, we would end up in state $q_{2}$.
II-B State Machines Considered in SEQUENT
State machines come in two flavors, namely deterministic or non-deterministic state machines [26]. SEQUENT solely considers deterministic state machines as they are easier to interpret; simpler to predict the next event, and verifying the same sequence of events would execute the same sequential behavior is more straightforward [15]. We can employ two approaches for learning a state machine from trace data: active and passive learning. We opt for the latter approach as it only requires us to collect traces from the target system $S$ instead of continuously interacting with $S$ to learn a state machine [2].
State-merging heuristic algorithm is a highly effective method for learning state machines in the passive-learning approach [13]. It first starts with a tree-like state machine (see Figure 2) known as the Prefix Tree Acceptor (PTA). It then iteratively compares pairs of states to identify states that could be merged. Merges are done based on the scores computed during the comparisons. The algorithm ends when it cannot find any new merges. The final output is an approximation of the smallest state machine consistent with the input data.
Figure 2: An example of a PTA.
III Assumptions of the Adversary
Our work is grounded in a set of assumptions we formulated concerning the objectives and capabilities of an adversary. We consider an adversary attacking an arbitrary target network $TN$, driven by various objectives such as infiltrating $TN$ to find machines containing sensitive data, disrupting normal operations by launching denial of service attacks against specific services, orchestrating a botnet, or infecting machines with malware. Our main experiments take the train and test data from the datasets to simulate an adversary attacking $TN$ without taking our defense into account. In our robustness experiments, we assume a more adept adversary that is aware of an ADS within $TN$ but unaware of its specifics. Their goal is to remain undetected until they have accomplished their objectives. To accomplish this, the adversary can monitor $TN$ to collect normal network behaviors and then attempt to bypass the ADS by generating behaviors resembling the ones that were collected. The adversary can only listen to the data that was sent but cannot delete any of the sent data.
IV Related Works
Traditional Methods. Several works have studied the efficacy of unsupervised learning for network anomaly detection. The primary focus has been on utilizing traditional unsupervised ML algorithms (e.g., Isolation Forest). Zoppi et al. conducted multiple studies evaluating the effectiveness of traditional unsupervised methods for anomaly detection on NetFlow data [31, 32]. In contrast to SEQUENT, these studies do not dynamically adapt the anomaly scores based on the events/traces seen at test time for anomaly detection.
State Machines. Numerous works have investigated the efficacy of state machines for anomaly detection. Pellegrino et al. presented BASTA, a system that learns probabilistic timed state machines from NetFlows and uses them as behavioral fingerprints for machines infected with botnet malware [24]. Alerts are raised when traces match behaviors captured in the fingerprints. Unlike SEQUENT, BASTA uses malicious data for learning as opposed to solely benign data.
Lin et al., Cao et al., and Matoušek et al. learned probabilistic state machines to detect anomalies in different systems[10, 16, 18]. These works raise alerts when a new trace deviates significantly from what was observed during training. The degree of deviation is quantified based on the likelihood of a trace. In contrast, SEQUENT quantifies the abnormality based on state visit frequency for a corresponding trace and compares this to what was observed during training.
Grov et al. presented a different approach for learning state machines to detect anomalies in NetFlows [12]. The approach learns state machines using solely traces of network protocols extracted from benign data. The degree of abnormality is based on the normalized edit distance of an arbitrary trace under scrutiny; the number of changes needed such that each trace event can be replayed in the state machine. In contrast to SEQUENT, this work does not utilize the frequency of the states to adapt the scores dynamically at test time. Furthermore, it solely considers network protocols for the learning, whereas SEQUENT considers other features.
Neural Networks. Deep Neural Networks (DNN) gained increasing popularity among researchers in recent years due to their ability to identify complex hidden patterns that traditional models may struggle to detect. Mirsky et al. introduced Kitsune, a state-of-the-art solution that utilizes an ensemble of Autoencoders (AE) to detect network anomalies in an unsupervised manner. It extracts various statistics from benign data and uses them for training. Kitsune then attempts to recreate instances it has seen during training by computing similar statistics for arbitrary new data. Alarms are raised for instances with large recreation errors[19]. Neural networks are notoriously known to be difficult to interpret, as it is not always clear what happens between the input and output layers. Nguyen et al. sought to address this issue with their framework, GEE. It uses a Variational Autoencoder (VAE) to detect anomalies and generates explanations by providing SHAP-like plots of the gradients computed for NetFlow features[23]. However, it is not clear how a security analyst should interpret these gradients. In contrast, SEQUENT learns an inherently interpretable model and generates root causes, which can be used to facilitate the understanding of (malicious) sequential behaviors. Root causes are linked back to concrete instances producing the anomalous behaviors.
V SEQUENT
V-A Intuition Behind SEQUENT
SEQUENT employs a novel approach for computing the anomaly score of an arbitrary trace under scrutiny. Unlike the measures of abnormality presented in prior works [24, 18, 16, 12, 10], SEQUENT leverages the frequencies of the state visits in a state machine to compute an adaptive anomaly score. The underlying principle of SEQUENT is that behaviors occurring significantly more often than those observed during training are indications of anomalies. Traces under scrutiny are replayed on a state machine to determine how well they fit within the learned model. Each trace visits a set of states, and the frequency of visits to each state is compared to the expected frequency observed during training. When the frequency of visits to a state is significantly higher than its expected frequency, it signals that unusual behavior is occurring in this state. The more significant the increase in visits for this state, the more likely it is to show anomalous behavior. SEQUENT employs this rationale to dynamically adapt its scores for arbitrary traces seen at test time. SEQUENT’s intuition will also work for (anomalous) behaviors never seen before during training, as these will visit unusual paths during the replay.
V-B SEQUENT’s Anomaly Detection Pipeline
Figure 3 depicts SEQUENT’s anomaly detection pipeline. It offers a high-level overview showing the flow of processing NetFlows, learning of a state machine, and finally generating the results. Each step is described in the following paragraphs.
Figure 3: SEQUENT’s anomaly detection pipeline.
Processing NetFlows. NetFlow needs to be discretized before learning a state machine, given that it contains both continuous and discrete feature data. Each flow needs to be transformed into an event to generate traces. The discretization process is not strictly necessary for the learning of a state machine, but using the raw NetFlow feature data (specifically continuous data) will lead to the combinatorial explosion of states. Discretization enables us to lower the number of possible combinations in the feature data, reducing the number of states used in the state machine. To discretize NetFlows, we develop an encoding algorithm inspired by the GloVe algorithm [25]. Our encoding algorithm computes co-occurrences between values for a given flow feature $f$ and stores them in a matrix. Each row denotes a unique value $v$ of $f$ and how often it occurs with another value $v^{\prime}$. Feature values are discretized by first clustering the rows using KMeans and then using cluster assignments as their discretization. We make our encoding algorithm available on Github [3]. To transform a flow into an event, we concatenate the discretized feature values with an underscore symbol to form a single event symbol. For example, if three features are used, we get the following event symbol: $DFV1\_DFV2\_DFV3$, where $DFV$ is the abbreviation for discretized feature value. In the case of one feature (e.g., bytes), we directly use its discretized values. To convert events into traces, we apply a sliding window of length $N$ to generate traces. These are then subsequently used to learn a state machine. Traces that do not have a length of $N$ are excluded from learning as they do not provide sufficient information.
Learning of a State Machine. SEQUENT utilizes FlexFringe [28] to learn a state machine from NetFlow traces. This framework offers, out of the box, various heuristics for learning a state machine. Additionally, it enables its users to learn other types of models e.g., Markov Chains. Furthermore, FlexFringe automatically creates traces from CSV files provided as input data. The user has to specify which features to use to create events, the length of the sliding window, and how the data should be sorted (e.g., by connections) to create the traces. We opt for FlexFringe based on these aforementioned features.
Computing Rolling Anomaly Score. Following the intuition described in Subsection V-A, we construct a formula for computing a rolling anomaly score for a state $q_{i}\in Q$ visited by an arbitrary trace $t$:
anomaly score (qi)=observations of qi so $\textit{anomaly score $(q_{i})$}=\frac{\sum\textit{observations of $q_{i}$ so far}}{E}$
where $E$ is defined as:
$E=\frac{\sum\ \textit{observations of $q_{i}$ in training traces}}{\sum_{i=0}^{n}\ \textit{observations of $q_{i}$ in training data}}\cdot\textit{UC}$
Then, UC is defined as the unique number of states seen so far in new traces:
$UC=\sum_{j=0}^{m}|StateSequence(t_{m})|$
and $StateSequence$ is a function that returns a set of states visited by a trace $t$:
$StateSequence(t)\rightarrow\{(q_{1},q_{2},...,q_{n})\},\textit{where }q_{i}\in Q$
UC only concerns the set of states learned during training, i.e., no new states are learned at test time. As $t$ can visit a particular state $q_{i}$ multiple times (e.g., if $q_{i}$ is part of a loop within the state machine model), this could skew the observations of $q_{i}$ (thus also the abnormality of $q_{i}$) for subsequent traces. To reduce this impact, we take the logarithm of the anomaly score computed for $q_{i}$, resulting in the following formula:
anomaly score ($\textit{anomaly score $(q_{i})$}=\log\left(\frac{\sum\ \textit{observations of$q_{i}$ so far}}{E}\right)$
This formula enables us to change the level of abnormality for a state $q_{i}$ based on how frequently it is visited by each $t$ seen at test time. The anomaly score for $t$ is the sum of the anomaly scores computed for the set of states it has visited.
Computing Root Causes. An advantage of using a state machine to model sequential behavior is the ability to follow and analyze the path that an arbitrary trace $t$ has taken within the model. We leverage this capability to compute a root-cause symbol for $t$. We do so by finding the state with the largest rolling anomaly score visited by $t$ and using the state identifier as the root-cause symbol. It denotes the state that causes the largest anomalous activity in $t$. As an example, let $t=[a,a,b,f,a,b,c]$ be the trace under scrutiny, $q=[1,4,7,8,4,4,4]$ be the sequence of states visited by $t$ and $a=[0.19,0.20,0.18,0.16,0.50,0.60,0.65]$ be the rolling anomaly scores computed for $q$. State 4 has the largest score (0.65) and thus ”4” is the root-cause symbol for $t$.
Grouping Anomalies. A key benefit of SEQUENT is the ability to create groups of anomalies exhibiting similar network behavior. SEQUENT leverages the root-cause symbols to group anomalies. The rationale is that traces sharing the same root-cause symbol denote behaviors exhibiting a notable level of anomalous activity in the same state. As traces signify the paths taken within the state machine, and the anomaly score of a state increases based on its visit frequency, traces sharing the same root-cause symbol will frequently visit this particular state. These traces will partly consist of the same events.
Linking Root Cause Back to Data. Another benefit of SEQUENT is its ability to link root causes to the corresponding flow, producing a notable level of anomalous activity. SEQUENT achieves this by utilizing the insights produced by FlexFringe after making predictions on a set of arbitrary traces. For each prediction of a trace $t$, FlexFringe reports details on which lines of the CSV file were used for the creation of $t$. Specifically, the start and end points are reported. SEQUENT utilizes this information to pinpoint the exact flow within the CSV file, producing a notable level of anomalous activity in $t$ and links this to the root-cause symbol. This enables us to collect concrete instances of NetFlows for anomalies that are grouped under the same root-cause symbol.
VI Experiment Setup
VI-A Dataset Selection
To showcase SEQUENT’s effectiveness in identifying network anomalies in an unsupervised setting, we have selected three publicly accessible datasets. These datasets comprise malicious NetFlows originating from a wide range of attacks. This allows us to assess SEQUENT’s ability to identify various types of attacks. Furthermore, the varied nature of these datasets enables us to evaluate SEQUENT’s generalizability across different network architectures.
The first dataset consists of NetFlows collected from a Kubernetes cluster[9] and used in the study conducted by Cao et al. [10]. It contains benign flows generated by real-life users and malicious flows generated by three attack scenarios. The three scenarios are constructed using a specialized threat matrix of the MITRE ATT&CK framework, designed explicitly for Kubernetes [10]. Each attack is fragmented into various steps that an adversary has to execute (e.g., initial access, execution, discovery, etc.). The three attack scenarios cover attacks such as (1) exploiting a vulnerability to discover secrets and using them to do privilege escalation, (2) executing a denial of service (DoS) attack, and (3) exploiting a vulnerability that allows the deployment of malware.
The second dataset is a well-known NetFlow dataset consisting of a large collection of realistic botnet network traffic mixed with normal and background traffic. Specifically, there are thirteen captures of various botnet samples. Each comprises various actions and protocols exhibited by the deployed malware [11]. The captures cover a wide range of attacks, e.g., distributed DoS, port scans, Click-Fraud, and sending spam. The benign and background flows were collected from normal machines deployed on a university network, and malicious flows were collected from virtual machines that were running the different botnets. Each virtual machine was bridged onto the university network to collect flows for the final dataset.
The third and final dataset consists of 100 GB of NetFlows collected from a Spanish ISP. The benign flow data was collected from real-life clients of the ISP, and the malicious flow data was collected from five machines that were used to launch the attacks. The attacks were not launched against any clients but rather five machines that were set up as victim machines [17]. The dataset mostly covers various types of DoS and scanning attacks. Furthermore, it also covers attacks in which spam emails are sent and botnet traffic. The botnet traffic was collected from the CTU-13 dataset and was modified before they were inserted into this dataset. The modifications consist of matching the IP addresses and timestamps of the flows to the ones used in their setup.
VI-B Creating Train/Test Splits
We outline the procedure for creating the train/test split for each dataset. Before the split is created, NetFlows are first sorted by timestamp and then by connections. We believe connection data has less noise than data sorted by timestamps, and this would help with learning more accurate models.
The AssureMOSS dataset was collected over a two-hour period, during which no attack scenarios were initiated within the first hour. We leverage the first hour as our train split and the second hour as our test split. This provides us with 120 thousand flows for training and 275 thousand flows for testing.
As for the CTU-13 dataset, we follow the split presented by its authors [11]: scenarios 3,4,5,7,10,11,12 and 13 for training and scenarios 1,2,6,8,9 for test. The training scenarios contain malicious flows, so we only use benign flows for training. Moreover, using all data from each scenario will lead to an overly large training and test set. To prevent this, we sample a subset of flows from each scenario, resulting in 640 thousand flows for training and 522 thousand flows for test.
For the UGR-16 dataset, we follow the split presented by its authors [17]: data from March to June are used as training data, and data from July and August are used as test data. As this dataset contains, by far, the largest amount of NetFlow data, we use the same sampling technique as with CTU-13 to create our train and test set. This yields a total of 600 thousand flows as training data and 664 thousand flows as test data.
VI-C Methods for Comparisons
We collected methods previously evaluated on the selected datasets to assess how SEQUENT compares to existing unsupervised and state-of-the-art methods. Additionally, we included various methods that serve as baselines.
From Literature. The AssureMOSS dataset was recently published by Cao et al. and remains relatively unexplored in the context of evaluating other unsupervised anomaly detection methods. Cao et al. conducted a preliminary evaluation comparing their approach against Isolation Forest and Local Outlier Factor [10].
Regarding the CTU-13 dataset, Zoppi et al. evaluated various unsupervised methods for identifying network anomalies [31, 32]. Specifically, the authors evaluated the efficacy of Isolation Forest and Fast Angle-Based Outlier Detection (FastABOD). Additionally, Grov et al. evaluated the efficacy of Markov Chains in detecting botnets [12]. Pellegrino et al. also conducted a study in which they evaluated the efficacy of (probabilistic) state machines and Bigram models (which are essentially Markov Chains) in detecting botnets [24].
For the UGR-16 dataset, Zoppi et al. evaluated the use of two different methods to detect network attacks [31, 32]. However, due to their inability to handle the size of our train/test splits, we dropped them for comparison. Nguyen et al. utilized this dataset to evaluate the efficacy of GEE. Similar to Kitsune [19], they use a state-of-the-art deep learning technique to detect anomalies. However, GEE explicitly trains a single VAE instead of an ensemble of AE.
We select the following from the compiled list of methods to compare against SEQUENT: Isolation Forest, FastABOD, Markov Chain, (probabilistic) state machines, and GEE. As for the implementation of the selected methods, we use Isolation Forest from the scikit-learn library [8], FastABOD from the PyOD library [30], Flexfringe to learn a Markov Chain and (probabilistic) state machine, and the implementation of GEE written by Wong Mun Hou [29]. For each method, except for the Markov Chain, we use the same parameters reported by their authors to train a model on the corresponding dataset. In cases where no parameter information is provided, we make our best effort to train a model by following the reported training steps. We chose to learn a Markov Chain using FlexFringe as we employ a distinct set of features, and learning a Markov Chain follows a standard procedure.
Additional Methods. We included additional methods beyond the aforementioned ones to compare against SEQUENT. One of these methods is the straightforward Boxplot Method presented by Arp et al.  [6], as it employs the frequency of data to detect anomalies. For our experiment, we wrote our implementation of the Boxplot Method following the description provided by the authors. Zoppi et al. also used Histogram-Based Outlier Detection (HBOD) to detect network anomalies. Although not applied to any of the selected datasets, it utilizes the frequency of bins to detect anomalies. This serves as a nice alternative baseline in our comparison. We use PyOD’s implementation of HBOD. Finally, we included Kitsune as it is a state-of-the-art unsupervised ADS. We use the implementation provided by the authors [20]. As it was designed to run at the packet level, we have to make minimal changes for it to work at the flow level. We recognize that this might negatively impact its performance, but we believe that it is beneficial to have this as a baseline for assessing SEQUENT’s usefulness.
VI-D Feature Selection & FlexFringe Setup
As selected datasets do not report the same flow features, we are required to select a subset of features shared among the selected dataset. The features shared between the selected datasets are timestamp, duration, protocol, number of bytes sent, number of packets sent, source and destination IP addresses, and source and destination ports. From this subset of features, we select the duration, protocol, number of bytes sent, and the number of packets sent to generate traces for SEQUENT. Port numbers are not considered within SEQUENT as they are spurious features within the CTU-13 dataset [22]. We discretize all the selected features except for protocol, as it is already a categorical feature.
We configure FlexFringe to sort the flows based on the source and destination IP pairs as we want to learn a state machine from connection data. Furthermore, we configure FlexFringe to use a sliding window of length 10 for the creation of traces. By default, FlexFringe stops replaying a trace if it reaches a state and no outgoing transition exists for the next event in the trace. We, therefore, configure FlexFringe to allow the replay to reset back to the root state and continue with the remaining events. This enables us to compute the visit frequencies of the states for the complete trace. Our FlexFringe configuration settings are available on our repository [4].
VII Evaluation Results
(a) AssureMOSS dataset.
(b) CTU-13 dataset.
(c) UGR-16 dataset.
Figure 4: ROC Curves of each model across the three selected datasets.
VII-A Effectiveness of SEQUENT
We report SEQUENT’s efficacy in detecting various network anomalies on the selected datasets. We use the Area Under Curve (AUC) score as the comparison metric between the various types of models. We deem this a suitable metric for comparison as each method generates anomaly scores for different traces, and we are uncertain about the optimal classification threshold that would yield the most favorable outcomes. Besides the AUC scores, we also report the Receiver Operator Characteristics (ROC) curves computed for each model. The AUC scores and ROC curves on each dataset are reported in Table I and Figure 4, respectively. These are average performance results computed from 10 independent runs. All models were trained on the same features considered by SEQUENT. We also include models trained on the original feature set used in the original study. These are marked with the $\dagger$-symbol in the table. Furthermore, FF-MC denotes the Markov Chain learned using the FlexFringe framework and FF-SM the (probabilistic) state machine. The reported results demonstrate that SEQUENT consistently outperforms all other models in detecting network anomalies on each respective dataset. Moreover, the results also suggest that SEQUENT is generalizable over different network architectures.
Model AssureMOSS CTU-13 UGR-16 Isolation Forest 0.140 0.473 0.426 Isolation Forest - 0.672 - FastABOD 0.229 0.492 0.576 FastABOD - 0.688 - HBOD 0.195 0.357 0.326 HBOD - 0.589 - Boxplot Method 0.436 0.410 0.738 FF-MC 0.067 0.232 0.392 GEE 0.335 0.473 0.505 GEE - - 0.683 Kitsune 0.371 0.5 0.621 FF-SM 0.041 0.142 0.257 SEQUENT 0.978 0.874 0.933
TABLE I: Average AUC of each model across all datasets.
Figure 5: Simplified state machine model learned from CTU-13 dataset.
Figure 6: Zoomed-in state machine of Figure 5, showing the top three anomalous states (47, 62 and 31) presented in Table II.
VII-B Example Analysis Using Root Causes
One of SEQUENT’s advantages is the ability to group anomalies using the generated root-cause symbols. This enables an analyst to reduce the time spent analyzing false positives by filtering out symbols deemed to be false positives. We provide an illustrative example to demonstrate how these symbols can aid with the analysis of anomalies. Let’s say that an analyst has set an arbitrary threshold on the (rolling) anomaly scores to select ”interesting” alerts for further analysis. The analyst then groups the selected traces based on their root-cause symbols and computes the size of each group. Table II presents the top 3 largest groups of anomalies computed from the CTU-13 dataset. Additionally, Figure 5 illustrates the simplified state machine learned by SEQUENT. The analyst decides to first analyze group ”62” by investigating the corresponding (starting) flows collected and linked by SEQUENT (see Table III). The most anomalous flows are shown at the top. SEQUENT deems these flows to be anomalous due to the significant increase in network behavior that coincidentally correlates with the use of port 25, which is used for Simple Mail Transfer Protocol (SMTP). The analyst deems these flows to be suspicious due to the destination port used in the communication and marks all traces with the root-cause symbol ”62” as malicious. Examining the actual labels of these flows, they correspond to the malicious traffic produced by the Neris botnet. This malware specifically sends spam mail [11], which explains the use of the SMTP port (25).
Root-Cause Symbol Size 62 18646 47 3182 31 2556
TABLE II: Anomaly group sizes computed using root-cause symbols generated for the CTU-13 dataset
src_ip dst_ip sport dport num_bytes num_packets
147.32.84.165 66.94.237.64 3271 25 186 3
147.32.84.165 98.139.175.225 3539 25 186 3
147.32.84.165 205.188.146.194 2474 25 186 3
147.32.84.165 76.96.30.116 3268 25 186 3
147.32.84.165 216.157.130.15 1490 25 186 3
147.32.84.165 98.137.54.237 3630 25 186 3
147.32.84.165 64.12.138.161 1789 25 186 3
147.32.84.165 209.85.143.27 3328 25 186 3
147.32.84.165 205.188.190.2 3857 25 186 3
147.32.84.165 98.139.175.225 3508 25 186 3
TABLE III: Simplified flows linked to the top 10 anomalous traces for root-cause symbol ”62”.
Following the same approach, the analyst now investigates group ”47” using the linked flows depicted in Table IV. SEQUENT does not know the actual nature behind these flows, but due to the significant increase in this type of network behavior, it deems these flows to be anomalous. Upon inspection, the analyst observes that these are correlated with the use of port 53, which is typically used for domain name resolution (DNS). This significant increase could be related to a DNS flood attack, or it could simply be that this type of network behavior was rarely seen during training. The analyst deems these flows unrelated to an attack and marks all traces grouped under the symbol ”47” as false positives. By investigating only a subset of the linked flows, the analyst can filter out 3182 false positive alerts and reduce the time spent analyzing every alert. Furthermore, the analyst can visualize the part of the model causing the most anomalous activity (see Figure 6) to understand the sequential behavior exhibited by the most anomalous groups of traces.
src_ip dst_ip sport dport num_bytes num_packets
147.32.84.138 147.32.80.9 45170 53 214 2
147.32.84.138 147.32.80.9 57876 53 214 2
147.32.84.138 147.32.80.9 50385 53 214 2
147.32.84.138 147.32.80.9 45753 53 214 2
147.32.84.138 147.32.80.9 34683 53 214 2
147.32.84.138 147.32.80.9 42564 53 214 2
147.32.84.138 147.32.80.9 56081 53 214 2
147.32.84.138 147.32.80.9 58016 53 214 2
147.32.84.138 147.32.80.9 53136 53 214 2
147.32.84.138 147.32.80.9 40272 53 214 2
TABLE IV: Simplified flows linked to the top 10 anomalous traces for root-cause symbol ”47”.
VII-C SEQUENT’s Robustness
An essential characteristic of an effective ADS is its ability to withstand attacks launched by adversaries seeking to evade detection. With various works showing that ML models are susceptible to adversarial attacks [5, 7, 27], it is imperative to evaluate SEQUENT’s efficacy against (potential) adversarial attacks. To so do, we constructed four distinct types of adversarial attacks following the adversary’s capabilities presented in Section III. We strictly consider attacks that an adversary could launch at test time, and we do not consider adversarial attacks such as poisoning of training data. An adversary is only allowed to adapt the features considered by SEQUENT for anomaly detection, and they are allowed to collect benign flows from the targeted network before the start of an attack. The constructed adversarial attacks are defined as follows:
  • Padding Attack: The adversary is allowed to mimic normal behavior by adapting each feature value to its nearest value within the collected data.
  • Random Replacement Attack: The adversary mimics normal behavior by replacing their flows with random flows selected from the collected data.
  • Window Replacement Attack: The adversary mimics normal behavior by replacing each sliding window (10 flows) of their flows with a randomly selected window of benign flows. Each window is selected at most once.
  • Frequency Replacement Attack: The adversary mimics normal behavior by finding frequently occurring flows (at least 100 times) from the collected data and replaces their flows with randomly selected frequently occurring flows. The intuition behind this attack is that an adversary would be able to blend in well with the normal flows if they mimic the frequently occurring communication behavior.
Tables V presents SEQUENT’s robustness against the aforementioned adversarial attacks on each selected dataset. The results are the average AUC scores across 10 independent runs. Based on the reported results, we see that SEQUENT achieved the best performance against two types of adversarial attacks (Padding Attack and Random Replacement Attack) compared to all other models. This can be explained by the nature of these two attacks; when an adversary tries to mimic the normal network behavior by padding or doing replacements on their flows, the traces will resemble the traces observed at training. However, this also changes the sequential behavior exhibited in the traces, which also results in changes in the state visit frequencies during the replay. SEQUENT detects this unusual increase in visit frequencies in the states and will raise an alarm. Existing methods trained on the raw feature value will not detect this deviation as these values are similar to what was observed during training. Probabilistic models such as Markov Chains and probabilistic state machines also suffer from this issue as the adversarial traces are similar to what was observed during training. The models will assign higher probabilities to these traces, thus resulting in poor detection performance.
Another interesting observation is that SEQUENT’s performance deteriorates significantly on the AssureMOSS and UGR-16 datasets but not on the CTU-13 dataset when an adversary launches a Window Replacement Attack. This decline in performance is expected since all malicious windows are randomly replaced by benign windows. Unlike the other three attacks, flows or feature values are not randomly replaced; instead, the flows and their sequential behaviors resemble the ones seen during training. As each window is allowed to be selected at most once, the sequential behavior exhibited in the (copied) window occurs at a slightly higher frequency, though not significantly more than observed during training. The complexity behind this attack makes it challenging for SEQUENT to detect the adversary. However, SEQUENT exhibits robustness against this specific attack on the CTU-13 dataset. We speculate that this result is influenced by the type of attack captured in the dataset.
One discussion point that we would like to initiate concerning three of the constructed adversarial attacks (in which replacement is involved) is whether an adversary could still accomplish their objectives by producing exact copies of the normal traffic data to avoid detection by an ADS. We argue that by producing literal copies and even going to the extent of recreating the sequential behaviors of benign flows, the adversary is essentially replaying collected network traffic data. No system, not even the analyst doing manual inspection of each flow or packet, would be able to explain how the data produced by an adversary differs from the normal traffic data as it is a literal copy.
Model Padding Random Window Frequency
AssureMOSS
Isolation Forest 0.162 0.185 0.431 0.155
FastABOD 0.222 0.220 0.177 0.031
HBOD 0.195 0.226 0.404 0.195
GEE 0.576 0.441 0.427 0.916
Kitsune 0.859 0.862 0.573 0.835
FF-MC 0.014 0.062 0.0319 0.985
FF-SM 0.007 0.038 0.278 0.985
Boxplot 0.435 0.465 0.502 0.441
SEQUENT 0.991 0.903 0.155 0.771
CTU-13
Isolation Forest 0464 0.484 0.515 0.478
FastABOD 0.492 0.475 0.436 0.290
HBOD 0.357 0.369 0.392 0.497
GEE 0.470 0.372 0.522 0.490
Kitsune 0.5 0.5 0.5 0.5
FF-MC 0.231 0.2 0.283 0.298
FF-SM 0.142 0.091 0.323 0.441
Boxplot 0.5 0.5 0.5 0.5
SEQUENT 0.970 0.966 0.861 0.974
UGR-16
Isolation Forest 0.431 0.429 0.480 0.470
FastABOD 0.588 0.566 0.468 0.336
HBOD 0.327 0.333 0.364 0.373
GEE 0.115 0.122 0.113 0.052
Kitsune 0.621 0.621 0.621 0.621
FF-MC 0.323 0.308 0.491 0.640
FF-SM 0.174 0.166 0.518 0.811
Boxplot 0.740 0.758 0.426 0.742
SEQUENT 0.983 0.919 0.407 0.927
TABLE V: The performance (average AUC) of each model against the four adversarial attacks.
VIII Conclusion & Future Work
We introduced SEQUENT, a framework that employs state machines to detect various types of network anomalies. SEQUENT leverages the structure of a state machine to collect state visit frequencies and subsequently uses them to continually adapt its scoring for arbitrary traces seen at test time. We observed that SEQUENT achieves a better overall performance in detecting various types of network anomalies, even outperforming existing state-of-the-art unsupervised anomaly detection methods. Furthermore, we demonstrated how SEQUENT computes root causes for the detected anomalies and uses them to group and rank anomalies. Moreover, SEQUENT links root causes back to the concrete NetFlow instances. The root causes, and the linked instances enable an analyst to quickly filter out a large number of false alerts, reducing the time spent on analyzing them. Finally, we evaluated SEQUENT’s robustness against four types of adversarial attacks. Reported results show that SEQUENT achieved the best performance in two of the adversarial attacks.
One potential future research direction is to investigate how the intuition underlying SEQUENT’s scoring method can be applied to other existing models, such as isolation forests or neural networks. It is not trivial how the internal structure of these models can be employed to continuously adapt its anomaly scores for new data seen at test time.
Another research direction is to apply SEQUENT to software log data. We believe that SEQUENT might also be effective in detecting anomalies in software logs, and it would be interesting to compare SEQUENT against existing anomaly detection methods developed for software log data.
Acknowledgments
The actual acknowledgment text is removed for authors’ anonymity and will be added after the review process. The current text is used as a placeholder so that we can take the required space into account later.